Matrix Methods/Linear Algebra: Please show all work and justify the answer! Just need Part C, the null Space and Part D please.

3 -6 9 0 1 -2 0 -6 3. Let A= 2 -4 7 2 The RREF of Aiso 0 1 2 3 -6 6 -6 0 0 0 (a) (6 points) Find a basis for Col A, the colum(d) (2 points) What is the dimension of the null space of A?
Answer #1

ANSWER

G AX=0 A=li -2 0-6 12 0 0 0 - 6 XI -90 O 01 AX=0 > X2 D. 0 OOD X3 *4) N-202-6%q=0 13 +2144=0 13=-904 10,=QW₂+644 Null(A) quat